Question 7 (3 points) If f(x) = 2x - 7, find: a. f(3x +2) b.f-?x)
Answer #1

Question 7 f(x) = 2x - 7 @). f(3x+2) = 2 (3x + 2) - 7 6X + 4-7 6x - 3 And (b). f(x) - 2x - 7 Let y 2X - 7 27 = y + 7 J x こ >